OpenStudy (anonymous):

Determine Laplace transform of f(t)=10 sin(2t-30degrees). 1.F(s)=20/(s+2)^2+4 2.20cos30deg./s^2+4 3.-5s+17.32/s^2+4 4.8.66/s^2+4

OpenStudy (anonymous):

whts laplace?

OpenStudy (anonymous):

its gauassians function

OpenStudy (anonymous):

gauassians?

OpenStudy (anonymous):

Lt(f(t) 10sin(2t-30 deg) F(s) =(20/(s+2^2)+4)(s) =-10(s-2sqrt3)F(s)/2s^2+8 =20/(s+2^2)+4/s

OpenStudy (anonymous):

\[L \left\{ \sin(2t - \pi/6) \right\} = \int\limits_{0}^{\infty}e ^{-st}\sin(2t - \pi/6) dt.\] Let's call that integral L and apply integration by parts twice to set up a linear equation in L. Here I let u be e^-st and dv be the trig function dt. I come up with \[L = \sqrt{3}/4 - s/8 - (s ^{2}/4)L.\]

OpenStudy (anonymous):

Solving for L , then multiplying by 10, gives us \[10 L = (10 \sqrt{3} - 5s)/(s ^{2} + 4).\] for the answer, at least according to my calculations.